OpenStudy (anonymous):

First one to answer gets a medal!!! What fraction is equal to the decimal number 0.4.

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

No not correct

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

0.4=4/10=2/5

OpenStudy (anonymous):

use your calculator to find the value of 2/5

OpenStudy (anonymous):

4/10

OpenStudy (anonymous):

you will find it to be equal to 0.4

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

4/10 IS 2/5 duh!

OpenStudy (anonymous):

4/10 is indeed 2/5

OpenStudy (anonymous):

\[\frac{4}{10}=\frac{2*2}{2*5}=\frac{2}{5}\]

OpenStudy (anonymous):

1/2.5

OpenStudy (anonymous):

IT IS 2/5 >_>

OpenStudy (anonymous):

derp

OpenStudy (anonymous):

4/10<=>2/5 what's the problem?